You are analyzing two companies that manufacture electronic toys—IntelliGames Inc. and BrainGames Inc. IntelliGames was launched eight years ago, whereas BrainGames is a relatively new company that has only been in operation for the past two years. However, both companies have an equal market share with sales of $100,000 each. You’ve gathered up company data to compare IntelliGames and BrainGames.

For the same period, the average sales for industry competitors was $255,000. As an analyst, you want to make comments on the expected performance of these two companies in the coming year. You’ve collected data from the companies’ financial statements, and the information follows:

Data Collected (in dollars)

IntelliGames Inc.

BrainGames Inc.

Industry Average

Accounts receivables $2,700 $3,900 $2,875
Net fixed assets 55,000 80,000 216,750
Total assets 95,000 125,000 234,600

Using the preceding information, complete the following statements.

1. A high or low days sales outstanding, or average collection period, represents an efficient credit and collection policy. Between the two companies,  IntelliGames or BrainGames    is collecting cash from its customers faster than IntelliGames or BrainGames ; but both companies are collecting their receivables less quickly than the industry average.

2. BrainGames Inc.’s fixed-asset turnover ratio is higher or lower than that of IntelliGames Inc. This could be because BrainGames is a relatively new company, such that the acquisition costs and book values of its fixed assets is greater or lower than the acquisition costs and book values of IntelliGames’s net fixed assets.

3. IntelliGames’ total asset turnover ratio is 1.05x or 0.80 , which is lower or higher than the industry’s average total asset turnover ratio. In general, a higher total asset turnover ratio indicates greater efficiency.
Answer #1

Days Sales outstanding = 365*accounts receivables/sales

IntelliGames = 365*2700/100,000 = 9.855 days

BrainGames = 365*3900/100,000 = 14.235 days

Industry Average = 365*2875/255,000 = 4.115 days

Fixed Assets turnover ratio = sales/fixed assets

Intelligames = 100,000/55,000 = 1.82 times

BrainGames = 100,000/80,000 = 1.25 times

Total Asset Turnover ratio = Sales/Total Assets

IntelliGames = 100,000/95,000 = 1.05

BrainGames = 100,000/125,000 = 0.8

Industry Average = 255,000/234,600 = 1.09

1.A lower/less DSO represents efficient credit policy. IntelliGames is collecting cash faster BrainGames.

2.BrainGames’s fixed asset turnover ratio is lower than Intelligames. Acquisition cost is GREATER

3.Intelligames total asset turnover ratio is 1.05, which is LOWER